A paint manufacturer desires a cylindrical steel drum to hold 50 gallons of roof paint.For convenience in handling,it is found necessary to limit the inside diameter of 2 and one-half feet.Find the height of the drum desired.
Answer by checkley77(12844)

50 gal=50/7.48=6.68 ft^3
6.68=h*area
6.68=h*pir^2
6.68=h*3.14(2.5/2)^2
6.68=h*3.14(1.25)^2
6.68=h*3.14*1.5625
6.68=h*4.906
h=6.68/4.906
h=1.36 feet tall for the drum.